# and Tight DP Auditing

### v2

@@ 7.3 Recommendations
Auditing Details. As discussed in Section 6.1, one not-so-uncommon limitation in DP auditing research is that key details of the auditing methodology are spread across various sections of the paper and may not be explicitly stated or left unclear. We suggest that researchers working on DP auditing explicitly state the DP guarantee audited, along with the adjacency notion, sub-sampling scheme, threat model, and confidence interval considered.
For instance, they could adopt an “Auditing Card” to be added in the appendix and/or in code repositories, similar to the “Model Cards” released for models published on HuggingFace.444See https://huggingface.co/docs/hub/en/model-cards. We include a draft template in Figure 3; future work could conduct a user study with researchers and practitioners working on DP auditing and explore the viability and opportunity of this approach.
Tighter -DP or -GDP guarantees. As also discussed by Gomez et al. [48], we also advocate for the standardized release and auditing of tighter -DP or -GDP guarantees. For example, when auditing the Gaussian mechanism, future audits could estimate and compare the empirical value to the theoretical -GDP guarantee, instead of converting both to a ()-DP value with an arbitrarily chosen . For mechanisms without a -GDP guarantee, the full -DP curve could be estimated, and the -divergence [58, 48] between the estimated and theoretical curve could be used as a “goodness of fit” metric.
@@ 7.4 Open Research Questions
In conclusion, our systematization effort highlights the presence of at least three open research questions.
@@
RQ2: Are end-to-end and tight audits possible? One of the most elusive combinations of desired properties is end-to-end-ness and tightness. While this is possible for simple mechanisms, for complex algorithms like DP-SGD, very little work has come close. Based on our systematic analysis, we believe this question can be further broken down into investigating slightly weaker threat models, namely, (Final Model, Gradient Canary) and (Intermediate Models, Sample Canary) before addressing the end-to-end threat model, i.e., (Final Model, Sample Canary). Cebere et al. [19] audit DP-SGD in (Final Model, Gradient Canary) but fail to achieve tightness, suggesting that tighter theoretical analysis might be necessary, which is in itself an open research question [5, 85]. Hence, we call for future work to focus on the (Final Model, Gradient Canary) and (Intermediate Models, Sample Canary) threat models.
@@ Appendix A Scopus Search Query
Appendix B Prior Work on DP auditing

### v3

@@ 7.2 Recommendations
Auditing Details. As discussed in Section 6.1, one not-so-uncommon limitation in DP auditing research is that key details of the auditing methodology are spread across various sections of the paper and may not be explicitly stated or left unclear. We suggest that researchers explicitly state the DP guarantee audited, along with the adjacency notion, sub-sampling scheme, threat model, and confidence interval considered.
For instance, they could adopt an “Auditing Card” to be added in the appendix and/or in code repositories, similar to the “Model Cards” released for models published on HuggingFace.666See https://huggingface.co/docs/hub/en/model-cards. We include a draft template in Figure 3; future work could conduct a user study with researchers and practitioners working on DP auditing and explore the viability of this approach.
Tighter -DP or -GDP guarantees. As also discussed by Gomez et al. [50], we advocate for the standardized release and auditing of tighter -DP or -GDP guarantees. For example, when auditing the Gaussian mechanism, future audits could estimate and compare the empirical value to the theoretical -GDP guarantee, instead of converting both to a ()-DP value with an arbitrarily chosen . For mechanisms without a -GDP guarantee, the full -DP curve could be estimated, and the -divergence [60, 50] between the estimated and theoretical curve could be used as a “goodness of fit” metric.
@@ 7.3 Open Research Questions
In conclusion, our systematization effort highlights the presence of at least four open research questions.
@@
RQ2: Are end-to-end and tight audits possible? One of the most elusive combinations of desired properties is end-to-end-ness and tightness. While this is possible for simple mechanisms, for complex algorithms like DP-SGD, very little work has come close. Based on our systematic analysis, we believe this question can be further broken down into investigating slightly weaker threat models, namely, (Final Model, Gradient Canary) and (Intermediate Models, Sample Canary) before addressing the end-to-end threat model, i.e., (Final Model, Sample Canary).
Cebere et al. [21] audit DP-SGD in (Final Model, Gradient Canary) but is not tight, suggesting that tighter theoretical analysis might be necessary, which is in itself an open research question [6, 89]. Hence, we call for future work to focus on the (Final Model, Gradient Canary) and (Intermediate Models, Sample Canary) threat models.
@@
RQ4: How do we make audits more robust? Although we have mainly focused on the three desired properties mentioned previously, recent research [111] has suggested that other properties might be important to achieve in DP auditing as well. Specifically, Wang et al. [111] find that in some cases, it may be possible for malicious actors to pass DP auditing techniques with overstated privacy guarantees, reducing the reliability and robustness of audit techniques.
On the other hand, even though theoretical DP guarantees are all-encompassing, holding for all adversaries and inputs, DP audits are fundamentally empirical and can only test a finite subset of them. As a result, current DP auditing techniques only focus on a single class of attacks, which makes it challenging for them to identify bugs that are only observable under specific classes of attacks. This also reduces the robustness of current DP auditing techniques and makes them brittle in practice. Overall, we find that the robustness of DP audits is another important property of DP auditing that is currently understudied and deserving of the attention of privacy researchers.
Acknowledgements. This work has been supported by the National Science Scholarship (PhD) from the Agency for Science, Technology and Research, Singapore.
@@ Appendix A Scopus Search Query
Appendix B Prior Work
@@
| Koskela et al. (2025) [65] | Laplace Gaussian DP-SGD DP-SGD | HS Div. | – Poisson Poisson Poisson | – – MI MI | Density Est. | ✗ ✗ ✗ ✓ | ✓ ✓ ✓ ✗ | ✓ |